1. Evaluate the complex integral:   ∫C [zRe(z) − z¯Im(z)]dz, where C is the line segment joining −1 to i. (z¯ = z bar)

2. Evaluate the complex integral:   ∫ C [iz^2 − z − 3i]dz, where C is the quarter circle with centre the origin which joins −1 to i.
Answer #1

IM (Z) ☺sz Raz)- Jule)dz Recz) Z(t) = (t-1+ it 0<t<1 = Srce-+it) (t =– (=1)-it) ] (1+i) dt 12+ i f(t-1) – (t1)t tit?] (1+i)dt